OpenStudy (anonymous):

Show that the dual of the exclusive-OR is equal to its complement. Can anyone help me?

OpenStudy (anonymous):

ex-or = AB' + A'B ex-or dual = (A+B') (A'+B) therefore ex-or dual becomes AA' + AB + A'B' + BB' thus ex-or dual = AB + A'B' = ex-or' or complement of ex-or... hope you understood...:-)

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

you're welcome:)